Light with a wavelength of 493 nm in vacuum travels through a transparent substance.
(a) What is the wavelength, in nanometers, of this light in water, with a refractive index of 1.33?
(b) What is the wavelength, in nanometers, of this light in Plexiglass, with a refractive index of 1.48?
(c) What is the wavelength, in nanometers, of this light in flint glass, with a refractive index of 1.77?
1
Expert's answer
2020-11-20T07:13:07-0500

We can find the wavelength of light in the medium from the formula:


"\lambda=\dfrac{\lambda_0}{n},"

here, "\lambda_0=493\ nm" is the wavelength of light in vacuum, "\lambda" is the wavelength of light in the medium and "n" is the refractive index of medium.

a) For the case of water, we get:


"\lambda=\dfrac{493\cdot 10^{-9}\ m}{1.33}=371\cdot 10^{-9}\ m=371\ nm."

b) For the case of plexiglass, we get:


"\lambda=\dfrac{493\cdot 10^{-9}\ m}{1.48}=333\cdot 10^{-9}\ m=333\ nm."

c) For the case of flint glass, we get:


"\lambda=\dfrac{493\cdot 10^{-9}\ m}{1.77}=278.5\cdot 10^{-9}\ m=278.5\ nm."

Answer:

a) "\lambda=371\ nm."

b) "\lambda=333\ nm."

c) "\lambda=278.5\ nm."